f(x)=2.5x-10.5 substitute 2 for x
Answer by stanbon(75887)

f(x)=2.5x-10.5 substitute 2 for x
-----
f(2) = 2.5*2 - 10.5 = 5 - 10.5 = -5.5
---------------
Cheers,
Stan H.
------------